OpenStudy (anonymous):

How would I figure this out? The sum of two numbers is 50. The first number is 43 less than twice the second number. Write and solve a system of equations to find the two numbers.

OpenStudy (anonymous):

n1 + n2 = 50 (n2-43) + n2 = 50 2n2 - 43 = 50 2n2 =50 +43 n2 = 93/2 n2 = 46.5 and n1 = n2 - 43 = 46.5 - 43 = 3.5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sorry..., i wrong n1 + n2 = 50 ((2xn2)-43) + n2 = 50 3n2 - 43 = 50 3n2 =50 +43 n3 = 93/2 n3 = 31 and n1 = (2xn2) - 43 = (2x31) - 43 = 62 - 43 = 19 so n1 = 19 and n2 = 31 this is the real answer @kyrra

OpenStudy (anonymous):

its ok i see what you did thanks :)

OpenStudy (anonymous):

welcome :)